OpenStudy (anonymous):

Related Rates Question: Water is poured into a conical container at the rate of 10 cm3/sec. The cone points directly down, and it has a height of 30 cm and a base radius of 5 cm. How fast is the water level rising when the water is 5 cm deep (at its deepest point)?

OpenStudy (anonymous):



OpenStudy (anonymous):

V= pi/3 *r^2 *h = pi/3 (h/6)^2*h

OpenStudy (anonymous):

differentiate with respect to time... solve for dh/dt plug in the values given for dV/dt and h...

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

THank you!!!

OpenStudy (anonymous):

Sure!